A car rounds a curve bank at 30 degrees abd radius of 316m. find a. the necessary centripetal accelerationon this banked path so the car wont slip down or slip up the incline B. incline of the car?
Answer #1

let m is the mass of the car.

let a_rad is the necessary centripetal accelerationon

If theta is the angle of inclination,

Apply,

m*g*sin(theta) = m*a_rad*cos(theta)

m*g*tan(theta) = m*a_rad

==> a_rad = g*tan(theta)

= 9.8*tan(30)

= 5.66 m/s^2 <<<<<<-------------Answer